### v1

# Varying Newton constant,entropy and the black hole evaporation law

Abstract
In Einstein equations we represent the energy-momentum tensor as the one ($T^{\mu\nu}$ ) of an ideal fluid plus the cosmological term. We consider time-dependent Newton “constant” $G$, the cosmological term $\Lambda$ and non-conserved $T^{\mu\nu}$. The Bianchi identity imposes a relation between the energy-momentum (non)conservation and the variation of $G$ and $\Lambda$. If the energy-momentum $T^{\mu\nu}$ is conserved then both the Newton “constant” $G$ and the cosmological term $\Lambda$ either do not change or both must depend on time. If the energy-momentum $T^{\mu\nu}$ is not conserved then the Bianchi identity implies a relation between the energy-momentum and a variation in time of either $G$ or $\Lambda$ (or both). We apply thermodynamics in order to express the non-conservation of the energy-momentum of an ideal fluid by entropy and relate the time variations of $G$ and $\Lambda$ to a change of entropy. Using the relation between a varying Newton constant G and the black hole entropy we derive a modified formula for the Schwarzschild black hole evaporation (a slower evaporation). Its life time is $\frac{9}{5}$ times larger than the one for a constant $G$. The average black hole’s density remains constant when the black hole’s radius shrinks to zero .
1 Introduction
We consider Einstein equations
$G^{\mu\nu}=\tilde{T}^{\mu\nu},$ (1)
where $G^{\mu\nu}=R^{\mu\nu}-\frac{1}{2}g^{\mu\nu}R$ $\mu,\nu=0,1,2,3$ and $R^{\mu\nu}$ is the Ricci tensor(we use the conventions of [1]). The form of the energy-momentum tensor $\tilde{T}^{\mu\nu}$ comes from a phenomenological description of fields and matter (we call them fluids). Then, the equations of motion for the fluids can be derived [2][3] from the covariant conservation law of the energy-momentum. In fact, in a wider perspective, the covariant conservation law for $\tilde{T}^{\mu\nu}$ can be applied for a description of fluid dynamics of systems (e.g.,quark-gluon plasma [4])in a flat space with a geometry of a submanifold embedded in the Minkowski space-time .
The standard Einstein equations can be expressed in the form (velocity of light $c=1$)
$G^{\mu\nu}=8\pi GT^{\mu\nu}-g^{\mu\nu}\Lambda.$ (2)
[…]
In this paper we admit varying constants and a non-conservation of the energy-momentum $T^{\mu\nu}$ of an ideal fluid. The violation of the energy-momentum conservation can be a consequence of an incomplete theory ( e.g. some particles (fields) are not taken into account). Some approximations (e.g., diffusion approximation for a particle motion) directly imply a dissipative dynamics. The energy-momentum conservation can be violated if quantum phenomena are included in classical theory. We suggest that the quantum thermodynamic energy conservation (the first law of thermodynamics) can be identified with the conservation of $\tilde{T}^{\mu\nu}$ if we admit a time dependence of $G$ or $\Lambda$ (or both) in the classical Einstein equations (2).
In sec.2 we obtain Bianchi identities resulting from a conservation of $\tilde{T}^{\mu\nu}$(1) and non-conservation of $T^{\mu\nu}$ (2). We show that the conservation of $T^{\mu\nu}$ allows a variation in time of both $G$ and $\Lambda$ but $\partial_{t}G\simeq\partial_{t}\Lambda$. Such models have been discussed in [17][18][19][20].These authors show that a variation of $G$ leads to a variation of $\Lambda$ in an agreement with observations of the universe expansion. In sec.3 applying the thermodynamics we express the variation of cosmological constants by the entropy $S$. According to the first law of thermodynamics $\nabla_{\mu}T^{\mu\nu}$ is related to the heat flow. Then, owing to the second law of thermodynamics the heat flow should be expressed by the change of entropy $TdS$ (where $T$ is the temperature). From the Bianchi identities (applied to eq.(2)) and the first law of thermodynamics it follows that a change of entropy is impossible if $G$ and $\Lambda$ are constant. We discuss a change of $G$ in models with an entropy generated by quantum phenomena at the event horizon. We apply the relation between the entropy and a varying Newton constant $G$ to the Schwarzschild black hole evaporation. As a consequence we obtain a modified evaporation formula. We obtain a slower black hole evaporation. Its life time is $\frac{9}{5}$ times larger than the one for a constant $G$ (that the varying $G$ could change the evaporation law has been discussed in [11][12]). The black hole’s volume tends to zero, but the average energy density defined as a ratio of the mass by the volume does not change during the whole evaporation process ( in contradistinction with the case of a fixed Newton constant). The knowledge of an exact evaporation law may be important for recent searches of the primordial black hole explosions [21][22] raised by an observation of a giant emission of neutrino radiation in KM3NeT [23] (for a search of a final burst of radiation from primordial black holes see [24] and references cited there).
2 The Bianchi identity
Einstein equations (2) can be derived from the action
$W=\int dx\sqrt{-g}\Big(R-2\Lambda-8\pi GL\Big),$
where $g$ is the determinant of the metric,$L$ is the Lagrangian for “matter”fields or for an ideal fluid [25][26]. The Bianchi identity applied to eq.(1) leads to the equation
$\nabla_{\mu}\tilde{T}^{\mu}_{\nu}=0.$ (3)
It is understood that the solution of gravity in the presence of “fluids” means a solution of both eqs.(1) and (3).
The Bianchi identity (3) with $\tilde{T}^{\mu\nu}$ of eq.(2) gives an equation for the “fluids”
$8\pi\partial_{\mu}GT^{\mu}_{\nu}+8\pi G\nabla_{\mu}T^{\mu}_{\nu}-\partial_{\nu}\Lambda=0.$ (4)
We assume that the Riemannian manifold is globally hyperbolic . On a globally hyperbolic manifold one can introduce coordinates [27] such that the metric takes the form ($j,k=1,2,3$, we use the notation which will be applied in subsequent sections)
$dl^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}=-|g_{00}|dt^{2}+\gamma_{jk}dx^{j}dx^{k}.$ (5)
We assume that $G$ and $\Lambda$ depend only on time. Then, eq.(4) reads
$8\pi\partial_{t}GT^{0}_{\nu}+8\pi G\nabla_{\mu}T^{\mu}_{\nu}-\delta^{0}_{\nu}\partial_{t}\Lambda=0.$ (6)
At $\nu=0$ we have
$8\pi\partial_{t}GT^{0}_{0}+8\pi G\nabla_{\mu}T^{\mu}_{0}-\partial_{t}\Lambda=0.$ (7)
At the spatial index $j$
$\partial_{t}GT^{0}_{j}+G\nabla_{\mu}T^{\mu}_{j}=0.$ (8)
We assume the energy-momentum tensor $T^{\mu\nu}$ in the form of an ideal fluid
$T^{\mu\nu}=(\rho+p)u^{\mu}u^{\nu}+g^{\mu\nu}p,$ (9)
where $\rho$ is the energy density and $p$ is the pressure. The fluid velocity $u^{\mu}$ is normalized as $g_{\mu\nu}u^{\mu}u^{\nu}=-1$. For the ideal fluid we have
$u^{\nu}\nabla_{\mu}T^{\mu}_{\nu}=-(\rho+p)\nabla_{\mu}u^{\mu}-u^{\mu}\partial_{\mu}\rho.$ (10)
As an example of the metric (5) let us consider the FLRW metric describing an expanding isotropic homogeneous universe (with the curvature index k=1,0,1k=-1,0,1)
$dl^{2}=-dt^{2}+a^{2}\Big((1-kr^{2})^{-1}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\Big).$ (11)
In the metric (11) in the comoving frame $u=(1,{\bf 0})$ eq.(8) is satisfied if $p$ does not depend on spatial variables as $T^{0}_{j}=0$ and $\nabla_{j}T^{j}_{k}=\partial_{k}p=0$. The fluid velocity satisfies $\nabla_{\mu}u^{\mu}=(-g)^{-\frac{1}{2}}\partial_{\mu}((-g)^{\frac{1}{2}}u^{\mu})=3H$ with $H=a^{-1}\partial_{t}a$. Hence, if $\rho$ and $p$ depend only on time then from eqs.(7) and (9)
$8\pi\partial_{t}G\rho+8\pi G(\partial_{t}\rho+3H(\rho+p))+\partial_{t}\Lambda=0.$ (12)
If the energy-momentum tensor is conserved
$\partial_{t}\rho^{c}+3H(\rho^{c}+p^{c})=0$
then eq.(12) reads
$8\pi\partial_{t}G\rho^{c}=-\partial_{t}\Lambda.$ (13)
We obtain the standard formula for $\rho^{c}(a)$ from the energy-momentum conservation, but according to eq.(12) both $G$ and $\Lambda$ must depend on $a$. This is the case discussed by Gupta [17], Costa et al[18] and other authors [19] [20].
For the solution of eq.(13) it is useful to introduce the variable
$\tau=\ln a.$ (14)
We assume the linear equation of state $p^{c}=w^{c}\rho^{c}$ . From eq.(13) we can express $\Lambda$ by $G$ (we have $\rho^{c}=\rho^{c}_{0}a^{-3(1+w^{c})}$)
$\Lambda=\Lambda_{ex}-\rho^{c}_{0}\int_{\tau_{0}}^{\tau}8\pi\partial_{s}G\exp(-3(1+w^{c})s)ds,$ (15)
where $\Lambda_{ex}$ is an integration constant at $\tau_{0}$. We can interpret this formula as an appearance of a time-dependent cosmological term from a varying coupling constant [17][19].
In general, if the energy-momentum is not conserved, then we get from eq.(12)
$8\pi\partial_{t}(\rho G)+24\pi G\rho H(1+w)=-\partial_{t}\Lambda.$ (16)
By integration we can express $\Lambda$ in terms of $G\rho$ or $G\rho$ by $\Lambda$
$\begin{array}[]{l}\rho G=(\rho G)_{0}\exp(-3(1+w)(\tau-\tau_{0}))\cr-\frac{1}{8\pi}\int_{\tau_{0}}^{\tau}\exp(-3(1+w)(\tau-s))\partial_{s}\Lambda ds.\end{array}$ (17)
Let us note that choosing $\tau_{0}=$-$\infty$ (corresponding to the initial $a=0$) we get a solution with the initial $(\rho G)_{0}\exp(3(1+w)\tau_{0})=0$. If we assume $\Lambda=\Lambda_{ex}+\Lambda_{0}a^{\alpha}$, where $\Lambda_{0}$ is another constant, then in the Friedmann equation for the metric (11) we obtain from eq.(17) a modified matter distribution ( with $w_{\alpha}=-1-\frac{\alpha}{3}$; a phantom matter if $\alpha>0$ [6][28][29][30]).
3 Thermodynamics of the moving fluid
The first law of thermodynamics when applied to the energy-momentum $\tilde{T}^{\mu\nu}$ (1) means that the supply of heat $Q$ to the system in a volume $V$ yields a work $pdV$ done upon the system plus a change of its internal energy $E$
$dE+pdV=dQ=TdS.$ (18)
On the rhs of eq.(18) we have used the second law of thermodynamics (defining the entropy $S$) saying that there exists an integrating factor (the absolute temperature $T$) such that $dS=T^{-1}dQ$ depends only on the state of the system.
We are going to generalize eq.(18) so that it describes thermodynamics of an ideal fluid moving with a velocity $u^{\mu}$ in a metric more general than in eq.(11). Let us introduce the inverse temperature $\beta$ as a four-vector (for some recent discussions of an inverse four-temperature and the four-entropy, see [31][32][33] and references cited there)
$\beta^{\mu}=\beta u^{\mu}.$ (19)
The zeroth component
$\beta^{0}=\beta u^{0}=\beta|g_{00}|^{-\frac{1}{2}}(1+g_{jk}u^{j}u^{k})^{\frac{1}{2}}$ (20)
in the comoving frame $u=(u^{0},{\bf 0})$ is the Tolman-Ehrenfest equilibrium temperature [34], $T^{0}=T\sqrt{|g_{00}|}$, where $T=\beta^{-1}$ and $T^{0}=(\beta^{0})^{-1}$ (we choose units with the Boltzmann constant $k_{B}=1$). Then, we introduce the entropy density $\sigma^{\mu}$ as the four-vector
$\sigma^{\mu}=\sigma u^{\mu}.$ (21)
We express the first law of thermodynamics of a moving fluid as
$\beta^{\nu}\nabla_{\mu}T^{\mu}_{\nu}=-\nabla_{\mu}\sigma^{\mu}.$ (22)
We multiply the Bianchi identity (4) by $\beta^{\nu}$
$8\pi\partial_{\mu}G\beta^{\nu}T^{\mu}_{\nu}+8\pi G\beta^{\nu}\nabla_{\mu}T^{\mu}_{\nu}-\beta^{\nu}\partial_{\nu}\Lambda=0.$ (23)
Then, inserting (22) in eq.(23) we obtain
$8\pi\beta^{\nu}T^{\mu}_{\nu}\partial_{\mu}G-8\pi G\nabla_{\mu}\sigma^{\mu}-\beta^{\nu}\partial_{\nu}\Lambda=0.$ (24)
This is the basic equation which we intend to apply for further studies of gravitational systems with a cosmological horizon. In the example of the metric (11) we have (in the comoving coordinates)
$\begin{array}[]{l}\nabla_{\mu}\sigma^{\mu}=u^{\mu}\nabla_{\mu}\sigma+\sigma\nabla_{\mu}u^{\mu}\cr=u^{\mu}\nabla_{\mu}\sigma+3H\sigma=\partial_{t}\sigma+3H\sigma.\end{array}$ (25)
Hence, eq.(24) reads
$8\pi\beta\rho\partial_{t}G+\beta\partial_{t}\Lambda=-8\pi G(\partial_{t}\sigma+3H\sigma)$ (26)
or if we apply eq.(12)
$\partial_{t}\rho+3H{\cal G}=T\partial_{t}\sigma.$ (27)
where ${\cal G}=\rho+p-\sigma T$ is the density of the Gibbs potential. If we set $\rho=\frac{E}{V}$ and $\sigma=\frac{S}{V}$ in eq.(27) then we recover eq.(18).
4 Compact systems with an event horizon
Eq.(24) may have a wider application to space-time phenomena. We have in mind the star contraction [35] or universe evolution ( e.g.,for time-dependent metrics relevant for black holes in an expanding universe [36][37][38]) . In this section we apply an integrated form of eq.(24) to a compact object. We integrate eq.(24) with the measure $d{\bf x}\sqrt{-g}=d{\bf x}\sqrt{|g_{00}|}\sqrt{\gamma}$, where $\gamma$ is the determinant of the spatial metric in eq.(5) ( for a general metric a better way would be to do the space-time integral in the time-interval $(t,t+dt)$ as in [39])
$\int d{\bf x}\sqrt{g}\Big(\beta^{\nu}T^{\mu}_{\nu}\partial_{\mu}G-G\nabla_{\mu}\sigma^{\mu}-\frac{1}{8\pi}\beta^{\nu}\partial_{\nu}\Lambda\Big)=0.$ (28)
We assume that $G$ and $\Lambda$ depend only on time. In the comoving frame from eq.(20) $\beta^{0}=\beta|g_{00}|^{-\frac{1}{2}}$. Calculating the integral (28) of the energy-momentum of an ideal fluid and using $\nabla_{\mu}\sigma^{\mu}=(-g)^{-\frac{1}{2}}\partial_{\mu}((-g)^{\frac{1}{2}}\sigma^{\mu})$ we obtain
$\begin{array}[]{l}\partial_{t}G\int d{\bf x}\sqrt{\gamma}\beta\rho+G\partial_{t}\int d{\bf x}\sqrt{g}\sigma^{0}+G\int d{\bf x}\partial_{j}(\sqrt{\gamma}\sigma^{j})\cr+\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\beta\partial_{t}\Lambda=0.\end{array}$ (29)
We may skip the term $G\int d{\bf x}\partial_{j}(\sqrt{\gamma}\sigma^{j})$ as in the comoving frame $\sigma^{j}=\sigma u^{j}=0$. Denoting
$S=\int d{\bf x}\sqrt{\gamma}\sigma$ (30)
using $\sigma^{0}=u^{0}\sigma=|g_{00}|^{-\frac{1}{2}}\sigma$ and assuming that $\beta$ depends only on time we can rewrite eq.(29) as
$\begin{array}[]{l}\beta\partial_{t}G\int d{\bf x}\sqrt{\gamma}\rho+G\partial_{t}S+\frac{\beta}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda=0.\end{array}$ (31)
We define the mass
$M=\int d{\bf x}\sqrt{\gamma}\rho.$ (32)
We skip the time derivative of $\Lambda$ in eq.(31)for an application to the black hole. Then, eq.(31) can be rewritten as
$M\partial_{t}G=-GT\partial_{t}S.$ (33)
We intend to apply the formulae (2) and (24) to gravitational systems with an event horizon whose entropy and temperature has been determined by other methods [40][41][42]. In classical closed systems the total entropy does not change ( neither does it in quantum systems). In systems with an event horizon we may restrict ourselves to subsystems located inside or outside the horizon. Then, their entropy may depend on time. The corresponding calculations have been performed in quantum systems ( calculation of the entanglement entropy) [43][44][45]( see the reviews in [46][47]). We think that this entropy can be incorporated in eq.(33). Eq.(33) is a consequence of eq.(24). In eq.(24) we have replaced the covariant derivative of the energy-momentum tensor (the lhs of eq.(18)) appearing in the Bianchi identity (23) by the covariant derivative of the entropy density according to the first law of thermodynamics (the rhs of eq. (18)). The quantum results for the entropy and for the temperature (obtained by calculations of the partition function [43] or from the quantum gravity in the Euclidean formulation [42]) show that $S\simeq\frac{1}{\hbar}$ and $T$ is proportional to $\hbar$. Hence, their product on the rhs of eqs.(27) and (33) has a classical meaning.
The formula for the temperature of the Schwarzschild black hole of mass $M$ is [41]
$T_{BH}=\frac{\hbar}{8\pi GM}.$ (34)
It can be derived from the correlation functions of the quantum scalar fields on the Schwarzschild background or from the Rindler space approximation on the basis of the KMS periodicity argument. Another derivation is based on Euclidean quantum gravity in one-loop approximation [42]. The entropy of the Schwarzschild black hole can be calculated from the second law of thermodynamics $dM=T_{BH}dS_{BH}$ on the basis of eq.(34)
$S_{BH}=\frac{4\pi}{\hbar}M^{2}G$ (35)
(it is also obtained from a path integral as an entanglement entropy [42][43][44][45], in such a case it would appear in the semi-classical expansion of the effective action). We insert the black hole entropy (35) in eq.(33). Then, we obtain
$\frac{3}{2}\frac{d\ln G}{dt}=-\frac{d\ln M}{dt}.$ (36)
Hence,
$G=G_{0}M_{0}^{\frac{2}{3}}M^{-\frac{2}{3}}.$ (37)
The change of mass $M$ resulting from the black hole evaporation through an emission of massless particles is described by the formula [48] [49][50][51] (see also the reviews in [46][47])
$\frac{dM}{dt}=-\hbar\alpha M^{-2}G^{-2},$ (38)
where $\alpha$ is a number which can be obtained from Stefan-Boltzmann formula multiplied by the black hole surface and an estimate of emissivity [48] ( the formula does not apply if $T_{BH}$ is lower than the temperature of the environment and for a small time). Eqs.(38) and (33) express the thermodynamic conservation law. The black hole is releasing the energy and the entropy gained during a mass accretion in the process of black hole formation.
Eq.(38) has the solution
$M^{3}=M_{0}^{3}-3\hbar\alpha\int_{t_{0}}^{t}G^{-2}(s)ds,$ (39)
where $t_{0}$ is the time of the black hole formation. We can insert this $M$ in eq.(37) for $G$. Denote
$X=\int_{t_{0}}^{t}G^{-2}(s)ds.$ (40)
Then, eq.(37) can be rewritten as
$\frac{dX}{dt}=G_{0}^{-2}M_{0}^{-\frac{4}{3}}(M_{0}^{3}-3\hbar\alpha X)^{\frac{4}{9}}.$ (41)
The resulting equation for $G$ has the solution
$G=G_{0}\Big(1-\frac{5}{3}\hbar\alpha G_{0}^{-2}M_{0}^{-3}(t-t_{0})\Big)^{-\frac{2}{5}}.$ (42)
Inserting $G$ of eq.(42) into the formula (39) for the mass we obtain
$M=M_{0}\Big(1-\frac{5}{3}\hbar\alpha G_{0}^{-2}M_{0}^{-3}(t-t_{0})\Big)^{\frac{3}{5}}\equiv M_{0}Z^{\frac{3}{5}}.$ (43)
It follows that
$\begin{array}[]{l}T_{BH}=\frac{\hbar}{8\pi G_{0}M_{0}}Z^{-\frac{1}{5}}\end{array}$ (44)
The entropy tends to zero as $GM^{2}\simeq Z^{\frac{4}{5}}$, the luminosity $I$ of the black hole is proportional to $T_{BH}^{4}$ and to the area of the black hole ($A\simeq(GM)^{2}$ ). Hence, the luminosity behaves as
$I\simeq Z^{-\frac{2}{5}}$ (45)
tending to infinity at the end of the black hole’s life time with a speed lower than the one for a fixed Newton constant.
The black hole’s radius $2GM$ tends to zero as $Z^{\frac{1}{5}}$ but the average energy density defined as the mass divided by the volume remains constant as
$M\Big(\frac{4\pi}{3}(2GM)^{3}\Big)^{-1}=M_{0}\Big(\frac{4\pi}{3}(2G_{0}M_{0})^{3}\Big)^{-1}$ (46)
This is different than in the Hawking-Page model [52][51] when the averaged energy density tends to infinity at the end of black hole’s life time. The results (42)-(46) could be compared with observations of the gamma-ray bursts [24]. In particular, with the recent observations of the giant neutrino emissions [23] which by some authors [21][22] are interpreted as resulting from primordial black hole explosions.
The life time of the black hole according to eq.(43) is
$t_{f}=\frac{3}{5}G_{0}^{2}M_{0}^{3}\hbar^{-1}\alpha^{-1}.$ (47)
It is longer by a factor $\frac{9}{5}$ than the one of the Hawking-Page model (which would result from the assumption that $G=const$ in eq.(39)).
5 Summary and outlook
We have shown that the Bianchi identity of classical Einstein equations admits a violation of the ideal fluid energy-momentum conservation if the Newton “constant” G and the cosmological term depend on time. As a generalization of the first and the second laws of thermodynamics the covariant energy-momentum divergence can be expressed by the covariant divergence of the entropy density. Then, the entropy density enters the Bianchi identity. In this paper we have applied the formula for the change of entropy during the black hole evaporation to determine the time-dependence of $G$ in the Bianchi identity. As a consequence of a growing $G$ the black hole evaporation is slower and the average black hole density is constant in contradistinction to the one for a fixed $G$ (this could be subject to astronomical observations) . The time-dependence of $G$ is unavoidable if we wish to describe the non-conserved energy-momentum of an evaporating black hole by Einstein equations (2). The increasing $G$ could have a physical stabilizing origin preventing the appearance of an infinite density when the black hole’s radius tends to zero. Another problem deserving a study concerns the entropy of an expanding universe with a cosmological horizon. For such a system the varying cosmological constant in the Bianchi identity should be relevant.
Data availability statement No data were created or analysed in this study
Acknowledgement A part of the results of this paper is contained in JH’s bachelor thesis. JH thanks Dr. Aleksander Kozak for supervising the thesis.