### v2

# Varying Newton constant,entropy and the black hole evaporation law

Abstract
In Einstein equations we represent the energy-momentum tensor as the one ($T^{\mu\nu}$ ) of a fluid plus the cosmological term. We consider time-dependent Newton “constant” $G$, the cosmological term $\Lambda$ and non-conserved $T^{\mu\nu}$. The Bianchi identity imposes a relation between the energy-momentum (non)conservation and the variation of $G$ and $\Lambda$. The covariant divergence $\nabla_{\mu}T^{\mu\nu}$ can be related to the first law of thermodynamics. For compact systems of mass $M$ from the Bianchi identity we obtain a power-law relation $G\simeq M^{-\gamma}$ with $\gamma$ depending on pressure or entropy. We discuss radiation and a mass loss described by the Stefan-Boltzmann law. In this formula we insert an expression for the black hole area and its temperature $T$. The Bianchi identity together with a formula for temperature and entropy $S$ determines the index $\gamma$ in the relation between the Newton constant $G$ and the mass $M$. If the entropy $S$ is defined by the equation $dS=T^{-1}dM$ then $\gamma=1$ (the same as for zero pressure). If the formula of Bekenstein-Hawking entropy holds true for time-dependent $G$ then $\gamma=\frac{2}{3}$. We discuss consequences for the evaporation law of some modified expressions for the entropy appearing in effective models of gravity resulting from an interaction with matter fields. In particular, $\gamma=1$ leads to a constant evaporation temperature whereas $\gamma>1$ to a decreasing temperature and luminosity.
1 Introduction
We consider Einstein equations
$G^{\mu\nu}=\tilde{T}^{\mu\nu},$ (1)
where $G^{\mu\nu}=R^{\mu\nu}-\frac{1}{2}g^{\mu\nu}R$ $\mu,\nu=0,1,2,3$ and $R^{\mu\nu}$ is the Ricci tensor(we use the conventions of [1]). The form of the energy-momentum tensor $\tilde{T}^{\mu\nu}$ comes from a phenomenological description of fields and matter (we call them fluids). Then, the equations of motion for the fluids can be derived [2][3] from the covariant conservation law of the energy-momentum.
The standard Einstein equations can be expressed in the form (velocity of light $c=1$)
$G^{\mu\nu}=8\pi GT^{\mu\nu}-g^{\mu\nu}\Lambda.$ (2)
[…]
In this paper we admit varying constants and a non-conservation of the energy-momentum $T^{\mu\nu}$ . The violation of the energy-momentum conservation can be a consequence of an incomplete theory ( e.g. some particles (fields) are not taken into account). Some approximations (e.g., diffusion approximation for a particle motion) directly imply a dissipative dynamics. The energy-momentum conservation can be violated if quantum phenomena are included in classical theory. We suggest that the quantum thermodynamic energy conservation (the first law of thermodynamics) can be identified with the conservation of $\tilde{T}^{\mu\nu}$ if we admit a time dependence of $G$ or $\Lambda$ (or both) in the classical Einstein equations (2).
In sec.2 we obtain Bianchi identities resulting from a conservation of $\tilde{T}^{\mu\nu}$(1) and non-conservation of $T^{\mu\nu}$ (2). We show that the conservation of $T^{\mu\nu}$ allows a variation in time of both $G$ and $\Lambda$ but $\partial_{t}G\simeq\partial_{t}\Lambda$. Such models have been discussed in [19][20][21][22].These authors show that a variation of $G$ leads to a variation of $\Lambda$ in an agreement with observations of the universe expansion. In secs.3 and 4 we show that $\nabla_{\mu}T^{\mu\nu}$ can be related to the heat $Q$ according to the formula $dQ=TdS$ where $T$ is the temperature and $S$ is the entropy. We apply this formula to non-equilibrium systems with time-dependent variables. From the Bianchi identities (applied to eq.(2)) and the first law of thermodynamics it follows that a change of entropy is impossible if $G$ and $\Lambda$ are constant. We obtain an equation which relates the time derivatives of $G$ and the mass $M$ of a compact system. If the pressure in the system is zero then this equation implies that $GM$ is constant. In some systems the Bianchi identity leads to a power-law relation $G\simeq M^{-\gamma}$. In sec.5 we discuss a compact system with an event horizon and its mass decrease resulting from a radiation described by the Stefan-Boltzmann law. In such systems the temperature as well as the area of the horizon depend on $G$ and $M$. When the Bianchi identity determines the relation between $G$ and $M$ then the Stefan-Boltzmann law predicts an explicit dependence on time of both $G$ and $M$. We insert the Bekenstein-Hawking formula for the temperature in this radiation law. In a system with zero pressure as a consequence of a constant $GM$ we obtain constant temperature and constant luminosity during the whole evaporation process (no black hole explosion). We derive the same conclusion from the first law of thermodynamics if the entropy is determined by the equation $dS=T^{-1}dM$. We obtain a different result if we assume that Bekenstein-Hawking formula for the entropy applies without any change to models with a time-dependent $G$. In such a case the Bekenstein-Hawking temperature as well as the black hole luminosity tend to infinity at the end of the evaporation process but the mean black hole density remains constant. We discuss the evaporation formula for various possible $\gamma$ in some other models of gravity. In particular, $\gamma>1$ leads to a decreasing black hole temperature and a decreasing luminosity (that the varying $G$ could change the evaporation law has been discussed in [10][11]).
The compact systems with an event horizon could lead to useful tests (besides the ones discussed in [17][18]) of a varying Newton constant $G$ because the area and the temperature of the horizon depend on $G$. The knowledge of an exact evaporation law may be important for recent searches of the primordial black hole explosions [23][24] raised by an observation of a giant emission of neutrino radiation in KM3NeT [25] (for a search of radiation from primordial black holes see [26] and references cited there). A decreasing black hole temperature would be welcome in models of primordial black holes as a cold dark matter.
2 The Bianchi identity
Einstein equations (2) can be derived from the action
$W=\int dx\sqrt{-g}\Big(R-2\Lambda-8\pi GL\Big)+K_{GHY},$
where $K_{GHY}$ is the Gibbons-Hawking-York surface term [27][28], $x=(t,{\bf x})$, $g$ is the determinant of the metric,$L$ is the Lagrangian for “matter”fields or for a fluid [29][30]. The Bianchi identity applied to eq.(1) leads to the equation
$\nabla_{\mu}\tilde{T}^{\mu}_{\nu}=0.$ (3)
It is understood that the solution of gravity in the presence of “fluids” means a solution of both eqs.(1) and (3).
The Bianchi identity (3) with $\tilde{T}^{\mu\nu}$ of eq.(2) gives an equation for the “fluids”
$8\pi\partial_{\mu}GT^{\mu}_{\nu}+8\pi G\nabla_{\mu}T^{\mu}_{\nu}-\partial_{\nu}\Lambda=0.$ (4)
We assume that the Riemannian manifold is globally hyperbolic . On a globally hyperbolic manifold one can introduce coordinates [31] such that the metric takes the form ($j,k=1,2,3$, we use the notation which will be applied in subsequent sections)
$dl^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}=-|g_{00}|dt^{2}+\gamma_{jk}dx^{j}dx^{k}.$ (5)
We assume that $G$ and $\Lambda$ depend only on time. Then, eq.(4) reads
$8\pi\partial_{t}GT^{0}_{\nu}+8\pi G\nabla_{\mu}T^{\mu}_{\nu}-\delta^{0}_{\nu}\partial_{t}\Lambda=0.$ (6)
At $\nu=0$ we have
$8\pi\partial_{t}GT^{0}_{0}+8\pi G\nabla_{\mu}T^{\mu}_{0}-\partial_{t}\Lambda=0.$ (7)
At the spatial index $j$
$\partial_{t}GT^{0}_{j}+G\nabla_{\mu}T^{\mu}_{j}=0.$ (8)
We consider the energy-momentum tensor $T^{\mu\nu}$ of a fluid
$T^{\mu\nu}=(\rho+p)u^{\mu}u^{\nu}+g^{\mu\nu}p+\Sigma^{\mu\nu},$ (9)
where $\rho$ is the energy density, $p$ is the pressure and $\Sigma^{\mu\nu}$ describes a dissipation (“viscosity” ) of the fluid [32][33] . The fluid velocity $u^{\mu}$ is normalized as $g_{\mu\nu}u^{\mu}u^{\nu}=-1$. We have
$u^{\nu}\nabla_{\mu}T^{\mu}_{\nu}=-(\rho+p)\nabla_{\mu}u^{\mu}-u^{\mu}\partial_{\mu}\rho+u^{\nu}\nabla_{\mu}\Sigma^{\mu}_{\nu}.$ (10)
As an example of the metric (5) let us consider the FLRW solution describing an expanding isotropic homogeneous universe (with the curvature index k=1,0,1k=-1,0,1)
$dl^{2}=-dt^{2}+a^{2}\Big((1-kr^{2})^{-1}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})\Big).$ (11)
The metric (11) is a solution of Einstein equations with an ideal homogeneous fluid ($\Sigma=0$). In the metric (11) in the comoving frame $u=(1,{\bf 0})$ eq.(8) is satisfied if $p$ does not depend on spatial variables as $T^{0}_{j}=0$ and $\nabla_{j}T^{j}_{k}=\partial_{k}p=0$. The fluid velocity satisfies $\nabla_{\mu}u^{\mu}=(-g)^{-\frac{1}{2}}\partial_{\mu}((-g)^{\frac{1}{2}}u^{\mu})=3H$ with $H=a^{-1}\partial_{t}a$. Hence, if $\rho$ and $p$ depend only on time then from eqs.(7) and (10)
$8\pi\partial_{t}G\rho+8\pi G(\partial_{t}\rho+3H(\rho+p))+\partial_{t}\Lambda=0.$ (12)
A generalization of eqs.(11)-(12) to $\Sigma\neq 0$ is interesting because it may describe a pasting of FLWR solution with the Schwarzschild solution describing a black hole in an expanding metric [34][35][36][37].
If the energy-momentum tensor is conserved
$\partial_{t}\rho^{c}+3H(\rho^{c}+p^{c})=0$
then eq.(12) reads
$8\pi\partial_{t}G\rho^{c}=-\partial_{t}\Lambda.$ (13)
We obtain the standard formula for $\rho^{c}(a)$ from the energy-momentum conservation, but according to eq.(12) both $G$ and $\Lambda$ must depend on $a$. This is the case discussed by Gupta [19], Costa et al[20] and other authors [21] [22].
For the solution of eq.(13) it is useful to introduce the variable
$\tau=\ln a.$ (14)
We assume the linear equation of state $p^{c}=w^{c}\rho^{c}$ . From eq.(13) we can express $\Lambda$ by $G$ (we have $\rho^{c}=\rho^{c}_{0}a^{-3(1+w^{c})}$)
$\Lambda=\Lambda_{ex}-\rho^{c}_{0}\int_{\tau_{0}}^{\tau}8\pi\partial_{s}G\exp(-3(1+w^{c})s)ds,$ (15)
where $\Lambda_{ex}$ is an integration constant at $\tau_{0}$. We can interpret this formula as an appearance of a time-dependent cosmological term from a varying coupling constant [19][21].
In general, if the energy-momentum is not conserved, then we get from eq.(12)
$8\pi\partial_{t}(\rho G)+24\pi G\rho H(1+w)=-\partial_{t}\Lambda.$ (16)
By integration we can express $\Lambda$ in terms of $G\rho$ or $G\rho$ by $\Lambda$
$\begin{array}[]{l}\rho G=(\rho G)_{0}\exp(-3(1+w)(\tau-\tau_{0}))\cr-\frac{1}{8\pi}\int_{\tau_{0}}^{\tau}\exp(-3(1+w)(\tau-s))\partial_{s}\Lambda ds.\end{array}$ (17)
Let us note that choosing $\tau_{0}=$-$\infty$ (corresponding to the initial $a=0$) we get a solution with the initial $(\rho G)_{0}\exp(3(1+w)\tau_{0})=0$. If we assume $\Lambda=\Lambda_{ex}+\Lambda_{0}a^{\alpha}$, where $\Lambda_{0}$ is another constant, then in the Friedmann equation for the metric (11) we obtain from eq.(17) a modified matter distribution ( with $w_{\alpha}=-1-\frac{\alpha}{3}$; a phantom matter if $\alpha>0$ [5][38][39][40]).
3 Thermodynamics of the moving fluid
We multiply the Bianchi identity (4) by the velocity $u$, integrate with $\sqrt{-g}d{\bf x}$ and write it in a form showing a relation to the first law of thermodynamics
$\int d{\bf x}\sqrt{-g}\Big(u^{\mu}\partial_{\mu}(G\rho)+G(\rho+p)\nabla_{\mu}u^{\mu}+u^{\nu}\nabla_{\mu}\Sigma^{\mu}_{\nu})\Big)=-\frac{1}{8\pi}\int d{\bf x}\sqrt{-g}u^{\mu}\partial_{\mu}\Lambda,$ (18)
where eq.(10) has been applied. We express eq.(18) in the comoving frame in the metric (5) ( then $g_{00}u^{0}u^{0}=-1$ and $\sqrt{-g}=\sqrt{-g_{00}}\sqrt{\gamma}$, where $\gamma$ is the determinant of the spatial metric). We use
$\nabla_{\mu}u^{\mu}=(\sqrt{-g})^{-1}\partial_{\mu}(u^{\mu}\sqrt{-g})=(\sqrt{-g})^{-1}\partial_{t}\sqrt{\gamma}.$ (19)
Then, eq.(18) reads
$\begin{array}[]{l}\int d{\bf x}\sqrt{\gamma}(\partial_{t}G\rho+G\partial_{t}\rho)+G\int d{\bf x}(\rho+p)\partial_{t}\sqrt{\gamma}\cr=-\int d{\bf x}\sqrt{\gamma}\nabla_{t}\Sigma^{0}_{0}-\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda.\end{array}$ (20)
The first law of thermodynamics when applied to the energy-momentum $\tilde{T}^{\mu\nu}$ (1) means that the supply of heat $Q$ to the system in a volume $V$ yields a work $pdV$ done upon the system plus a change of its internal energy $E$
$dE+pdV=dQ=TdS.$ (21)
On the rhs of eq.(21) we have used the second law of thermodynamics (defining the entropy $S$) saying that there exists an integrating factor (the absolute temperature $T$) such that $dS=T^{-1}dQ$ depends only on the state of the system. Eq.(21) can be rewritten in the form
$d{\cal H}-Vdp=TdS,$ (22)
where ${\cal H}=E+pV$ is the enthalpy.
It can be seen that the lhs of eq.(20) coincides with the lhs of eq.(21) if we extend the meaning of the differentials of neighboring configurations to configurations at different times. The term $\Sigma$ related to the heat flow in phenomenological thermodynamics is usually also expressed by entropy. We shall assume that it is already contained in $T\partial_{t}S$ and ignore it. Note that if $p=0$ then eq.(20) could be interpreted as the formula (22) for enthalpy with $p$ replaced by $\frac{1}{8\pi}\Lambda$ in the formulation of the black hole pressure according to to ref.[41].
When we define the mass as
$M=\int d{\bf x}\sqrt{\gamma}\rho$ (23)
then eqs.(20)-(21) give
$\begin{array}[]{l}M\partial_{t}G+G(\partial_{t}M+\int d{\bf x}p\partial_{t}\sqrt{\gamma})=-\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda\end{array}$ (24)
or in terms of the entropy
$\begin{array}[]{l}M\partial_{t}G+GT\partial_{t}S=-\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda.\end{array}$ (25)
In eqs.(23) and (25) we introduced the volume (for the metric (5))
$dV=d{\bf x}\sqrt{\gamma}$ (26)
with $d\partial_{t}V=d{\bf x}\partial_{t}\sqrt{\gamma}$.
For $p=0$ eq.(24) takes a simple form which will be applied in sec.5
$\begin{array}[]{l}\partial_{t}(MG)=-\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda.\end{array}$ (27)
There is a long-standing debate over the notion of the energy, volume and the mass in the context of thermodynamics in general relativity [42] (for a recent discussion see [43]). In the classic paper [44] the mass has the ADM meaning [1] of the mass measured at infinity and evaluated on the basis of its gravitational effects. The mass (23) is the mass of the fluid. For relativistic stars eq.(23) agrees with Weinberg’s definition of the mass ( ref.[45], sec.11, eq.(11.1.23)) if $\rho$ is interpreted as the density of nucleons in the star. The mass (23) appears in the formulas for the neutron star. It leads to the relation $G\simeq M^{-\frac{2}{3}}$ ([45], sec.11)in the formula for the Chandrasekhar limt of the neutron star. According to ref.[18] it gives a bound on the variation of $G$. In geometries pasting the FLWR geometry (5) with the Schwarzschild geometry (as discussed in sec.5) these various definitions of the mass may be related.
The reason that we integrated the covariant equation (4) with $d{\bf x}\sqrt{-g}$ is that the space-time volume $dtd{\bf x}\sqrt{-g}$ in eq.(18) does not depend on the choice of coordinates. We could integrate in eq.(18) with $dtd{\bf x}\sqrt{-g}v$ with an arbitrary function $v$. Then, in the integrated form of eq.(24) ( independent of the choice of coordinates if $v$ is a scalar)
$\begin{array}[]{l}\int dt\Big(\tilde{M}\partial_{t}G+G\partial_{t}(\tilde{M}+G\int d{\bf x}vp\partial_{t}\sqrt{\gamma}\Big)\cr=-\frac{1}{8\pi}\int dt\int d{\bf x}\sqrt{\gamma}v\partial_{t}\Lambda\end{array}$
with $\tilde{M}=\int d{\bf x}\sqrt{\gamma}v\rho$. The only change would involve a change of measure $\sqrt{\gamma}\rightarrow v\sqrt{\gamma}$ in the definition of the mass in eq.(23) and the entropy in terms of an entropy density in eq.(25).
We suggest a generalization of eq.(20) to an explicitly covariant formula for relativistic fluids. Let us introduce the inverse temperature $\beta$ as a four-vector (for some recent discussions of an inverse four-temperature and the four-entropy, see [33][46][47] and references cited there)
$\beta^{\mu}=\beta u^{\mu}.$ (28)
The zeroth component
$\beta^{0}=\beta u^{0}=\beta|g_{00}|^{-\frac{1}{2}}(1+g_{jk}u^{j}u^{k})^{\frac{1}{2}}$ (29)
in the comoving frame $u=(u^{0},{\bf 0})$ is the Tolman-Ehrenfest equilibrium temperature [48], $T^{0}=T\sqrt{|g_{00}|}$, where $T=\beta^{-1}$ and $T^{0}=(\beta^{0})^{-1}$ (we choose units with the Boltzmann constant $k_{B}=1$). Then, we introduce the entropy density $\sigma^{\mu}$ as the four-vector
$\sigma^{\mu}=\sigma u^{\mu}.$ (30)
We express the first law of thermodynamics of a moving fluid as
$\beta^{\nu}\nabla_{\mu}T^{\mu}_{\nu}=-\nabla_{\mu}\sigma^{\mu}.$ (31)
We multiply the Bianchi identity (4) by $\beta^{\nu}$
$8\pi\partial_{\mu}G\beta^{\nu}T^{\mu}_{\nu}+8\pi G\beta^{\nu}\nabla_{\mu}T^{\mu}_{\nu}-\beta^{\nu}\partial_{\nu}\Lambda=0.$ (32)
Then, inserting (31) in eq.(32) we obtain
$8\pi\beta^{\nu}T^{\mu}_{\nu}\partial_{\mu}G-8\pi G\nabla_{\mu}\sigma^{\mu}-\beta^{\nu}\partial_{\nu}\Lambda=0.$ (33)
This is the basic equation which we intend to apply for further studies of gravitational systems with a cosmological horizon. We have $\nabla_{\mu}\sigma^{\mu}=u^{\mu}\nabla_{\mu}\sigma+\sigma\nabla_{\mu}u^{\mu}$. Hence, if we use (4) and (10) then eq.(31) can be rewritten in the form
$u^{\mu}\partial_{\mu}\rho+(\nabla_{\mu}u^{\mu}){\cal G}=\beta^{-1}u^{\mu}\partial_{\mu}\sigma-\beta^{-1}\beta^{\nu}\nabla_{\mu}\Sigma^{\mu}_{\nu}.$ (34)
where ${\cal G}=\rho+p-\beta^{-1}\sigma$ is the density of the Gibbs potential. As (in the comoving frame) $\nabla_{\mu}u^{\mu}=(\sqrt{\gamma})^{-1}\partial_{t}\sqrt{\gamma}\equiv V^{-1}\partial_{t}V$ then (ignoring $\Sigma$)
$\partial_{t}\rho+V^{-1}\partial_{t}V{\cal G}=T\partial_{t}\sigma.$ (35)
When we insert in eq.(35) $\rho=\frac{E}{V}$ and $\sigma=\frac{S}{V}$ then we recover eq.(21). We can also rewrite eq.(35) in the form
$\beta^{-1}\partial_{t}(V\sigma)=\partial_{t}(V\rho)+p\partial_{t}V$ (36)
leading to the correct formula for static systems $\sigma=\beta(\rho+p)$.
4 Compact systems
In this section we show that eq.(31) leads to the same conclusions as eqs.(24)-(25) and (27) of sec.3. We integrate eq.(33) with the measure $d{\bf x}\sqrt{-g}=d{\bf x}\sqrt{|g_{00}|}\sqrt{\gamma}$
$\int d{\bf x}\sqrt{-g}\Big(\beta^{\nu}T^{\mu}_{\nu}\partial_{\mu}G-G\nabla_{\mu}\sigma^{\mu}-\frac{1}{8\pi}\beta^{\nu}\partial_{\nu}\Lambda\Big)=0.$ (37)
We assume that $G$ and $\Lambda$ depend only on time. In the comoving frame $\beta^{0}=\beta|g_{00}|^{-\frac{1}{2}}$ from eq.(28). Calculating the integral (37) of the energy-momentum and using $\nabla_{\mu}\sigma^{\mu}=(-g)^{-\frac{1}{2}}\partial_{\mu}((-g)^{\frac{1}{2}}\sigma^{\mu})$ we obtain
$\begin{array}[]{l}\partial_{t}G\int d{\bf x}\sqrt{\gamma}\beta T^{0}_{0}+G\partial_{t}\int d{\bf x}\sqrt{-g}\sigma^{0}+G\int d{\bf x}\partial_{j}(\sqrt{-g}\sigma^{j})\cr+\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\beta\partial_{t}\Lambda=0.\end{array}$ (38)
We may skip the term $G\int d{\bf x}\partial_{j}(\sqrt{\gamma}\sigma^{j})$ as in the comoving frame $\sigma^{j}=\sigma u^{j}=0$. Denoting
$S=\int d{\bf x}\sqrt{\gamma}\sigma$ (39)
using $\sigma^{0}=u^{0}\sigma=|g_{00}|^{-\frac{1}{2}}\sigma$ and assuming that $\beta$ depends only on time we can rewrite eq.(37) in the form (25)
$\begin{array}[]{l}\beta\partial_{t}G\int d{\bf x}\sqrt{\gamma}T^{0}_{0}+G\partial_{t}S+\frac{\beta}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda=0.\end{array}$ (40)
When we skip the time derivative of $\Lambda$ in eq.(40)and denote $M=\int d{\bf x}\sqrt{\gamma}T^{0}_{0}$ (we admit that with $\Sigma^{0}_{0}\neq 0$ the spatial integrals over $\rho$ and $T^{0}_{0}$ may be different) then we obtain
$M\partial_{t}G=-GT\partial_{t}S.$ (41)
5 Mass loss in evaporation
In this section we discuss dynamics of a change of mass of a compact object. We assume that such dynamics can be described in the framework of classical Einstein gravity if $G$ and $\Lambda$ are allowed to change in time. We have two equations (24) and (25) for a relation between time derivatives of $G$ and $M$. We have in mind a dynamical system resulting from a star contraction which subsequently is decaying through an emission of radiation. An investigation of the dynamics of a star collapse to the black hole has been initiated by Oppenheimer and Snyder [49]. Then, Einstein and Straus [50] [51] studied a star (Schwarzschild metric) in an expanding universe by pasting the FLWR metric with the Schwarzschild solution. In this framework there is a lot of work recently on the formation of the black hole [52] [53], its evolution in the FLWR space-time [34][35][36][37] and final evaporation [54][55] (nonsingular black holes). In this section we apply eqs.(25) and (27) to a description of a decrease of mass as a result of an emission of radiation of a compact object (a black hole) with a hot event horizon.
A change of mass $M$ resulting from an emission of radiation from a surface can be obtained from the Stefan-Boltzmann law (we assume that the temperature $T$ is larger then the temperature of the environment and that the absorption from the environment can be neglected)
$\frac{dM}{dt}=-\lambda T^{4}A,$ (42)
where $\lambda$ is the Stefan-Boltzmann constant ($\lambda\simeq\hbar^{-3}$) and $A$ is the area of the surface. For a compact system with an event horizon we can obtain a system of equations (24)-(25) and (42) depending on $G$ and $M$ because the area and temperature are functions of $G$ and $M$. In eq.(24) we would need to know the pressure whereas in eq.(25) the entropy as functions of $G$ and $M$ in order to determine a relation between these variables. We consider a power-law Ansatz as a possible solution of eqs.(24)-(25)
$G=G_{0}M_{0}^{\gamma}M^{-\gamma}.$ (43)
It can be seen that if $\partial_{t}\Lambda=0$ and $p=0$ then from eq.(24) $\gamma=1$. In order to determine a relation between $G$ and $M$ from eq.(25) we need to know the temperature $T$ and entropy $S$. We consider a simple power-law Ansatz
$T=\theta G^{\alpha_{1}}M^{\alpha_{2}},$ (44)
$S=\kappa G^{\nu_{1}}M^{\nu_{2}}.$ (45)
We substitute it in eq.(25). Then, we obtain
$\gamma=\theta\kappa\nu_{2}(1+\theta\kappa\nu_{1})^{-1}$ (46)
with $\alpha_{1}=-\nu_{1}$$\alpha_{2}=1-\nu_{2}$. If in addition we demand the definition of temperature as
$\frac{1}{T}=\frac{\partial S}{\partial M}$ (47)
then we get the relation $\kappa\theta\nu_{2}=1$ and $\gamma=\nu_{2}(\nu_{1}+\nu_{2})^{-1}$.
As conjectured by Bekenstein [56] the entropy of the black hole is proportional to its area. The calculation of the temperature of the black hole in a system of quantum scalar particles on a Schwarzschild background [57] gave the result (the formula has been established for various quantum systems defined on the Schwarzschild background [58][59])
$T_{BH}=\frac{\hbar}{8\pi GM}$ (48)
Then, from the relations (44)-(47) we obtain the entropy
$S_{BH}=\frac{4\pi}{\hbar}M^{2}G$ (49)
If instead of (47) we use (48) and
$dS=T_{BH}^{-1}dM$ (50)
appearing in the black hole mechanics in [44] then the formula for the entropy is
$S=\frac{8\pi}{\hbar}\int GMdM.$ (51)
For a time-dependent $G$ eqs.(49) and (51) are different. We discuss all options (49), (51)and (44)-(45) leading to the relation (43) with different $\gamma$. The reason is that a time-dependent $G$ may appear in an effective field theory of various origin [60]: Higgs gauge theories with symmetry breaking [61][62],Branse-Dicke [10][11][63], $f(R)$ and higher derivatives gravities [64][63]. For a constant $G$ the entropy (49) has been derived as an entanglement entropy in various models of quantum field theory [58][60][65][66]. A time dependent temperature (possibly in eq.(51)) can appear if the black hole surface is time dependent ( a trapping horizon) in an expanding universe [67]. A renormalization of the Newton constant as in refs. [13][14][15][16] can lead to a modified formula for the entropy. A microscopic counting of states in loop quantum gravity [68] does not fix the coefficient for entropy in front of the area as well. We consider also the option
$S_{\kappa}=\frac{\kappa}{4G\hbar}A,$ (52)
where $\kappa=G_{eff}^{-1}G$ with an effective coupling $G_{eff}$ which can come from a calculation of the entropy from the partition function. With the entropy (52) we get the relation (43) with $\gamma=\frac{\kappa}{1+\frac{1}{2}\kappa}$ .
Summarizing, $\gamma=1$ for zero pressure in eq.(25) or the entropy (51), $\gamma=\frac{2}{3}$ for the Bekenstein-Hawking entropy (49). The remaining values of $\gamma\geq 0$ are rather hypothetical. However, they lead to interesting consequences. A constant $G$ (corresponding to $\gamma=0$ ) cannot describe evaporation by means of Einstein equations because owing to Bianchi identities decreasing $M$ (with constant $\Lambda$) requires a change of $G$.
We assume in all models $T\simeq(GM)^{-1}$ and $A\simeq(GM)^{2}$. When we insert the temperature and the area in eq.(42) then we obtain [69] [70][71][72] (see also the reviews in [73][74][75])
$\frac{dM}{dt}=-\hbar\alpha M^{-2}G^{-2},$ (53)
where $\alpha$ is a ($\hbar$ independent) number which can be obtained from the Stefan-Boltzmann constant and an estimate of emissivity [69].
If the pressure is zero in eq.(24) (as in the Oppenheimer-Snyder approximation) leading to eq.(27) and $\partial_{t}\Lambda$ is negligible in eq.(27) then (as $GM\simeq const$) we obtain from eq.(53)
$M_{t}=M_{0}-\hbar\alpha M_{0}^{-2}G_{0}^{-2}(t-t_{0}),$ (54)
where $t_{0}$ is the time of the black hole formation. We obtain the same result ($\gamma=1$) from eq.(25) with an application of the formulas (50)-(51) for the entropy.
With a varying $\Lambda$ and $\gamma=1$ we substitute from eq.(27) in eq.(53)
$\begin{array}[]{l}MG=G_{0}M_{0}-\int_{t_{0}}^{t}\frac{ds}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{s}\Lambda.\end{array}$ (55)
Then,
$M=M_{0}-\hbar\alpha\int_{t_{0}}^{t}d\tau\Big(G_{0}M_{0}-\int_{t_{0}}^{\tau}\frac{ds}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{s}\Lambda\Big)^{-2}$ (56)
Note that from eq.(55)it follows that if $\partial_{t}\Lambda=0$ then the temperature of the black hole as well as its luminosity are constant during the whole black hole’s life time (there is no black hole explosion, contrary to [72]). With the varying cosmological term as a consequence of eq.(55) the variation of the temperature and luminosity of the black hole depends on the sign of $\partial_{t}\Lambda$ .
If eqs.(48)-(49) are applied in eq.(25) then
$\begin{array}[]{l}\frac{3}{2}M\partial_{t}G+G\partial_{t}M+\frac{1}{8\pi}\int d{\bf x}\sqrt{\gamma}\partial_{t}\Lambda=0.\end{array}$ (57)
If $\partial_{t}\Lambda=0$ then from eq.(57) we obtain $\gamma=\frac{2}{3}$. Surprisingly, this is the same relation between the Newton constant and the mass which follows from the formula for the mass of the neutron star and its Chandrasekhar limit [45] (sec.11) According to ref.[18] it could serve as an estimate of the variation of the Newton constant.
Let us consider the general option (43) with $\gamma\neq 0$ . When we substitute the relation (43) between $G$ and $M$ in eq.(53) then we obtain the solution for $M$ ( when $\gamma\neq\frac{3}{2}$)
$M=M_{0}\Big(1-(3-2\gamma)\hbar\alpha G_{0}^{-2}M_{0}^{-3}(t-t_{0})\Big)^{\frac{1}{3-2\gamma}}\equiv M_{0}Z^{\frac{1}{3-2\gamma}},$ (58)
$G=G_{0}\Big(1-(3-2\gamma)\hbar\alpha G_{0}^{-2}M_{0}^{-3}(t-t_{0})\Big)^{-\frac{\gamma}{3-2\gamma}}\equiv G_{0}Z^{-\frac{\gamma}{3-2\gamma}}.$ (59)
When $\gamma=\frac{3}{2}$ then
$M=M_{0}\exp\Big(-\hbar\alpha G_{0}^{-2}M_{0}^{-3}(t-t_{0})\Big)$ (60)
It follows that ( for $\gamma\neq\frac{3}{2}$)
$\begin{array}[]{l}GM=G_{0}M_{0}Z^{\frac{1-\gamma}{3-2\gamma}}\end{array}$ (61)
$\begin{array}[]{l}GM^{2}=G_{0}M_{0}^{2}Z^{\frac{2-\gamma}{3-2\gamma}}\end{array}$ (62)
and
$\begin{array}[]{l}G^{-3}M^{-2}=G_{0}^{-3}M_{0}^{-2}Z^{\frac{3\gamma-2}{3-2\gamma}}\end{array}$ (63)
These equations have some consequences for the behaviour of the temperature, luminosity (L), entropy and the mean density (defined as the mass divided by the volume). Clearly, $M$ tends to zero for all $\gamma$ and $G$ tends to infinity. However, the dependence of the temperature, luminosity, entropy and the mean desity may substantially depend on $\gamma$. For the temperature and luminosity we have
$\gamma$ T,L time
$0<\gamma<1$ increasing to infinity finite
$\gamma=1$ constant all time
$1<\gamma<\frac{3}{2}$ decreasing to zero finite
$\gamma\geq\frac{3}{2}$ decreasing to zero infinite
For $\gamma\geq\frac{3}{2}$ the mass, the temperature and the luminosity tend to zero asymptotically in time.
Let us note that the entropy (62) is decreasing to zero if $\gamma<2$, is constant if $\gamma=2$ and is growing in time if $\gamma>2$. The mean density (63)is tending to infinity if $\gamma<\frac{2}{3}$, is constant for $\gamma=\frac{2}{3}$ and is tending to zero if $\gamma>\frac{2}{3}$ (this is different than in the Hawking-Page model [72]). The results could be compared with observations of the gamma-ray bursts [26]. In particular, with the recent observations of the giant neutrino emissions [25] which by some authors [23][24] are interpreted as resulting from primordial black hole explosions.
The life time of the black hole for $\gamma<\frac{3}{2}$ according to eq.(58) is
$t_{f}=\frac{1}{3-2\gamma}G_{0}^{2}M_{0}^{3}\hbar^{-1}\alpha^{-1}.$ (64)
It is infinite for $\gamma\geq\frac{3}{2}$ but the mass is decreasing asymptotically to zero.
6 Summary
We have shown that the Bianchi identity of classical Einstein equations admits a violation of the fluid energy-momentum conservation if the Newton “constant” G and the cosmological term depend on time. The covariant energy-momentum divergence of the fluid appearing in the Bianchi identity has the form of the heat flow. We define the mass of the fluid as a volume integral of its energy density. If the pressure of the system is zero then the Bianchi identity relates $\partial_{t}(GM)$ to the time derivative of the cosmological term. For a non-zero pressure an analogous formula involves the temperature and entropy. A radiation of the fluid (involving an energy loss) can be included in Einstein equations if the the Newton “constant” or the cosmological term depend on time. The radiation is described by the Stefan-Boltzmann law. The Stefan-Boltzmann formula depends on the temperature and the area of the emitting surface. When the collapse of the star is described by pasting of Schwarzschild solution with the FLWR solution for a fluid with a certain density, then the resulting black hole should have the mass $M$ related to this density. We use the mass $M$ for a determination of the temperature and the area. The black hole mechanics derived in classical Einstein gravity gives a relation between $G$, $M$, the entropy $S$ and temperature $T$. In some effective theories of gravity when $G$ is time-dependent this relation may be more involved. From the Bianchi identity we obtain an equation relating the time derivatives of $G$, $\Lambda$ and $M$ to the temperature $T$ and entropy $S$. We discussed various formulas for $T$ and $S$ some of them directly generalizing the Bekenstein-Hawking expression and some (for $\gamma\neq 1$ and $\gamma\neq\frac{2}{3}$) quite hypothetical. These assumptions give a relation $G\simeq M^{-\gamma}$. When the expression $G(M)$ is substituted in the Stefan-Boltzmann formula for the mass decrease then it gives an evaporation law without free parameters. Depending on the value of $\gamma$ the black hole temperature and luminosity may either tend to infinity at finite time (black hole explosion), be constant during the whole black hole life time or even tend to zero in finite or infinite time. The latter option is interesting for the proposals of the dark matter as consisting of primordial black holes. The study of the possible black holes evaporation laws should give useful hints for their search in astronomical observations.
Data availability statement No data were created or analysed in this study
Acknowledgement A part of the results of this paper is contained in JH’s bachelor thesis. JH thanks Dr. Aleksander Kozak for supervising the thesis. ZH thanks Hytham Alwrekat for discussions on some topics of this paper.